# Diffusing States and Matching Scores: A New Framework for Imitation Learning

### v1

@@ 4 Algorithm
The cost function in Eq. 4 involves the score function $\nabla\log q_{t}(s_{t}\,|\,s)$, which has a closed-form expression for most modern diffusion models including DDPM. Specifically, when the diffusion follows the OU process (Eq. 1), the score function takes the form $\nabla\log q_{t}(s_{t}\,|\,s)=(1-e^{-2t})^{-1}(se^{-t}-s_{t})$, which is exactly the noise added to the original sample $s$ during the diffusion process (recalling that for the OU process, $q_{t}(s_{t}\,|\,s)=\mathcal{N}(se^{-t},(1-e^{-2t})I)$). We denote this noise by $\epsilon$. Then, the cost function (Eq. 4) is equivalent to
$c^{(k)}(s):=\mathop{\mathbb{E}}_{t\sim U(T)}\mathop{\mathbb{E}}_{\epsilon\sim\mathcal{N}(0,I)}\left[\Big{\|}g^{e}(se^{-t}+\epsilon,t)-\epsilon\Big{\|}_{2}^{2}-\Big{\|}g^{(k)}(se^{-t}+\epsilon,t)-\epsilon\Big{\|}_{2}^{2}\right].$
This closely resembles the noise prediction in DDPM.
@@
It is worth to pause for a moment and contrast the procedure we derived above with more traditional inverse RL / adversarial imitation learning procedures. At the highest level, we do not explicitly train a discriminator and instead simply perform score matching on the expert and learner state distributions. Doing so only requires the standard regression-based score-matching optimization, which has shown to be more stable than adversarial training in certain domains. On a more theoretical level, the DS-Divergence we’re optimizing in our procedure is an upper-bound on the worst-case IPM (the total variation distance) one often considers in IRL (Abbeel & Ng, 2004; Ziebart et al., 2008; Swamy et al., 2021). On the positive side, minimizing this upper bound merely requires score matching rather than explicitly training a discriminator. Score matching gives one stronger theoretical guarantees w.r.t. second-order bounds than IPM-based methods, as we discuss in the following section. Intuitively, the benefit of matching the second moments and not just the first are performance bounds that scale with the variance of the data-generating process, which can be significantly tighter on problems where stochasticity is restricted to a subset of states (e.g. driving on highways, where merges are significantly more variable than actually being on the highway). On the negative side, for problems where we can a-priori specify a class of moments (e.g. autonomous driving, where it is relatively simple to write down the set of features a person could care about when driving), minimizing an upper bound can be needlessly complex. Thus, we believe that as the difficulty of specifying the class of reward functions grows, our method’s benefits over more traditional IPM-based IRL techniques grows as well.
@@ 5 Theoretical Results
    There exists $\epsilon_{\rm{score}}>0$ such that the estimator $g^{e}$ is accurate up to $\epsilon_{\rm{score}}$:
    $\mathop{\mathbb{E}}_{s\sim d^{\pi^{e}}}\mathop{\mathbb{E}}_{t\sim U(T)}\mathop{\mathbb{E}}_{s_{t}\sim q_{t}(\cdot\,|\,s)}\big{[}\big{\|}g^{e}(s_{t},t)-\nabla_{s_{t}}\log p^{\pi^{e}}_{t}(s_{t})\big{\|}_{2}^{2}\big{]}\leq\epsilon_{\rm{score}}^{2};$
@@
Item (a) is a standard assumption in diffusion process and score matching (Chen et al., 2022; 2023; Lee et al., 2022). When the function class $\mathcal{G}$ (hypothesis space) is finite, $\epsilon^{2}_{\rm{score}}$ typically scales at a rate of $\widetilde{O}\left(\ln(|\mathcal{G}|)/N\right)$ in terms of the number of samples $N$. For an infinite function class, more advanced bounds can be established (e.g., see Theorem 4.3 in Oko et al. (2023)). We emphasize that the discretization error arising from approximating the diffusion SDE using Markov chains is included in $\epsilon_{\rm{score}}$ as well. Item (b) is similar to Item (a) but for the sequence of score function estimators. It can be satisfied by applying Follow-the-Leader (FTL) since the loss functional for $g$ is a square loss. Other no-regret algorithms, such as Follow-the-regularized-leader (FTRL) or online gradient descent, can also be used. Under similar conditions as in Item (a), one typically achieves ${\rm{Regret}}(K)=\widetilde{O}(\sqrt{K})$. Item (c) can be satisfied as long as an efficient RL algorithm is applied at each iteration in 4. The rate of $\epsilon_{\rm{RL}}$ has been well-studied in the RL theory literature for various of MDPs. For instance, in tabular MDPs, one can achieve $\epsilon_{\rm{RL}}=\widetilde{O}(\sqrt{SA/M})$ where $S$ is the number of states, $A$ is the number of actions, and $M$ is the number of RL rollout samples. We emphasize that $\epsilon_{\rm{RL}}$ in this case does not scale with $H$ as it is defined as the upper bound on the averaged return instead of the cumulative return over $H$ steps.
@@
Next, we introduce the regularity assumptions on the diffusion process, which are standard in the literature on diffusion models
@@
The first condition ensures that the score function behaves well so we can transfer the DS divergence bound to a KL divergence bound. The second condition is needed for the exponential convergence of the forward diffusion process (i.e., the convergence to the standard Gaussian in terms of KL divergence). We note that the second condition is readily satisfied in certain simple cases such as when the state space is bounded.
Notably, if we disregard discretization errors, the Lipschitz constant will not appear in our theoretical results as long as it is finite, even if it is arbitrarily large. Similarly, the constant $m$ only appears within logarithmic factors so it can be exponentially large. However, when discretization errors are considered, both the Lipschitz constant and $m$ will appear in $\epsilon_{\rm{score}}$ and ${\rm{Regret}}(T)$ (see, e.g., Chen et al. (2022; 2023); Oko et al. (2023)).
Now we are ready to present our main theoretical result. Let $\pi^{(1:K)}$ denote the mixture policy of $\{\pi^{(1)},\dots,\pi^{(K)}\}$. Particularly, $\pi^{(1:K)}$ is executed by first choosing a policy $\pi^{(k)}$ uniformly at random and then executing the chosen policy. For any policy $\pi$, we define the variance of its return as ${\rm{Var}}^{\pi}\coloneqq{\rm{Variance}}(\sum_{h=1}^{H}c^{\star}(s_{h}))$. Below is our main theoretical result with proof provided in Appendix B.
@@
Under Assumptions 1 and 2, Algorithm 1 achieves the following instance-dependent bounds:
@@
where we define total error
$\epsilon\coloneqq\begin{cases}\epsilon^{2}_{\rm{score}}+\epsilon_{\rm{RL}}+{\rm{Regret}}(K)/K&\text{if }\pi^{e}\in\Pi,\\ \epsilon^{2}_{\rm{score}}+\epsilon_{\rm{RL}}+{\rm{Regret}}(K)/K+\epsilon_{\rm{mis}}&\text{if }\pi^{e}\not\in\Pi,\end{cases}$
and the misspecification $\epsilon_{\rm{mis}}\coloneqq\min_{\pi\in\Pi}\ell(\pi)$ where we recall that $\ell(\cdot)$ is defined in Eq. 2.
@@
It is known that the second-order bound is tighter and subsumes the first-order bound (Wang et al., 2024c). Importantly, previous results on second-order bounds for sequential decision-making have relied exclusively on the Maximum Likelihood Estimator (MLE) (Wang et al., 2023; 2024c; Foster et al., 2024; Wang et al., 2024b; d), which in general is not computationally tractable even for simple exponential family distributions (Pabbaraju et al., 2024). In this work, we show for the first time that a computationally tractable alternative — diffusion score matching, can achieve these instance-dependent bounds in the context of IRL. In particular, our second-order bounds (or the first-order bound) scale with the minimum of ${\rm{Var}}^{\pi^{e}}$ and ${\rm{Var}}^{\pi^{(1:K)}}$ (or minimum of $V^{\pi^{e}}$ and $V^{\pi^{(1:K)}}$). Hence, our bounds are sharper whenever one of the two is small. In contrast, bounds in these prior RL work do not scale with the min of variances associated with the learned policies and the comparator policy $\pi^{e}$. This subtlety is likely due to the difference between the IRL setting and the general RL setting instead of techniques in the analysis.
Improved sample efficiency with respect to the size of expert data.
Let’s now focus on sample complexity with respect to expert data, which is typically the most expensive data to collect in IL, to highlight the benefits of our method over previous approaches. For conciseness, we ignore all errors that are purely related to computation. Specifically, we set $\epsilon_{\text{RL}}\rightarrow 0$, as it only depends on the time spent to run the RL procedure, let the number of iterations $K\rightarrow\infty$ so $\text{Regret}(K)/K\rightarrow 0$, and assume $\pi^{e}\in\Pi$. Under these conditions, our bound solely depends on $\epsilon_{\text{score}}$. For simplicity, we further assume a finite score function class $\mathcal{G}$, neglect misspecification errors (i.e., $\nabla\log p^{\pi}_{t}(x)\in\mathcal{G}$ for all $\pi$), and assume the magnitude of the score loss is $O(1)$. Now standard regression analysis (Agarwal et al., 2019; Oko et al., 2023) yields $\epsilon_{\text{score}}=O(\sqrt{\log(|\mathcal{G}|)/N})$ where $N$ is the number of expert samples. Plugging this back into our second-order bound, we obtain
@@
5.1 Comparison to Discriminator-based (f-divergence and IPM) Methods
@@
Hence, score matching only needs relatively weaker score function class to represent complicated distributions to control certain powerful divergence including KL divergence and Hellinger distance. However, for $f$-divergence-based or IPM-based methods, we need a more expressive discriminator function class to make them serve as a tight lower bound of the divergences. This perhaps explains why in practice for generative models, score matching based approach in general outperforms more traditional GAN based approach. In our experiment, we conduct an ablation study in Section 6.3 to compare our approach to a GAN-based baseline under linear function approximation for score functions and discriminators, and we show that SMILING can still learn well even when the score functions are linear with respect to the state $s$.
@@
6 Experiments
We evaluate our method, SMILING, on a set of continuous control tasks from the Deepmind Control Suite (Tassa et al., 2018) and HumanoidBench (Sferrazza et al., 2024). These tasks vary in difficulty, as summarized in Table 1. We elaborate on the experimental setup below. While we will compare to multiple baselines including non-adversarial baselines, we focus on testing the our hypothesis that our score-matching based approach should outperform f-divergence based adversarial IL approach. Additional details can be found in Appendix C.
Table 1: Spread of evaluated environments. DMC denotes DeepMind Control Suite and HB denotes HumanoidBench.
Task Difficulty Benchmark
Ball in Cup Catch Easy DMC
Cheetah Run Medium DMC
Humanoid Walk Hard DMC
Humanoid Sit Hard HB
Humanoid Crawl Hard HB
Humanoid Pole Hard HB
Baselines. We include Discriminator Actor-Critic (DAC), IQ-Learn, and Behavior Cloning (BC) as baselines. Our DAC implementation is based on the original design by Kostrikov et al. (2018) that uses JS-divergence based GAN style objective. The implementation of IQ-Learn directly uses the official version. We tried several configurations provided in their implementation for the MuJoCo tasks and reported the best one. The BC performance is reported as the maximum average episode reward obtained across all checkpoints during training. We experimented with both squared loss and MLE for BC and found MLE is better. Consequently, all BC results below are using MLE.
RL. For DMC tasks, we use Soft Actor-Critic (SAC) (Haarnoja et al., 2018) as the RL algorithm for both DAC and our method, based on the implementation provided by Yarats & Kostrikov (2020). For HumanoidBench tasks, we adopt DreamerV3 (Hafner et al., 2023) using the implementation from the HumanoidBench repository (Sferrazza et al., 2024). Additionally, we maintain a separate “state buffer” that operates the same as the replay buffer but only stores states. This buffer is used to sample states to update the score function in our method and the discriminator in DAC.
Overall, we keep the implementation of SMILING and the baseline DAC as close as possible (e.g., using the same RL solver) so the only difference is different objective functions (i.e., SMILING learns score functions while DAC learns discriminators to approximate JS divergence). This allows us to demonstrate the benefits of our score-matching approach to the discriminator-based GAN-style approaches. For IQ-Learn, we simply use the official implementation.
Expert Policy and Demonstrations. We use the aforementioned RL algorithms (SAC and DreamerV3) to train expert policies. Expert demonstrations are collected by executing the expert policy in the environment for five trajectories. We set the horizon of all tasks to be 1K without early termination, and thus the resulting dataset consists of 5K steps per task.
6.1 Learning from State-only Demonstrations
Figure 3: Learning curves for learning from state-only data across five random seeds. The x-axis corresponds to the number of environment steps (also the number of policy updates). The y-axis is normalized such that the expert performance is one and the random policy is zero. Our method clearly outperforms all baselines in five tasks out of six, with a substantial performance gap against baselines in four of them.
Figure 3 illustrates the results across five control tasks when the expert demonstrations are state-only. We trained for 0.4M steps on Ball-in-cup-catch and for 3M steps on all other tasks. Our method matches expert performance and outperforms both DAC and IQ-Learn on five of them. Specifically, our method outperforms DAC and IQ-Learn by a large margin on four tasks: ball-in-cup-catch, humanoid-walk, humanoid-crawl, and humanoid-pole. For humanoid-sit, while none achieves expert performance, ours converges faster. For cheetah-run, although DAC initially shows faster convergence, it becomes unstable soon and oscillates significantly after 2M steps. In contrast, our algorithm remains much more stable and surpasses DAC in 3M steps. We provide video demos of SMILING and DAC on the humanoid-crawl and humanoid-pole tasks in Figure 1 to give a clear visual comparison of their performance differences.
We include BC in the plot as a reference, though it is not directly comparable since it uses state-action demonstrations. It serves as a performance upper bound for Behavioral Cloning from Observation (BCO) (Torabi et al., 2018), which infers the expert’s actions from state-only data before training. Nevertheless, we observe that both our method and DAC consistently outperform BC (and thus, BCO).
6.2 Learning from State-action Demonstrations
6.3 Expressiveness of Diffusion Models
In Section 5.1, we argued that score matching is more expressive than discriminator-based methods, which may explain why our approach generally outperforms DAC in previous experiments. To further support this argument, we conducted an ablation study on the cheetah-run task where our method previously showed a slower convergence compared to DAC. Now we remove the activation functions in both the discriminator of DAC and the score function of our method, making them purely linear, and then re-run the experiments. The results are in Figure 5. Different from the previous experiments where DAC outperformed our method, in this case, DAC’s performance degrades notably by showing slower convergence and significantly higher variance. In comparison, our method remains more effective and outperforms DAC clearly.
We attribute the performance degradation of DAC to the limited expressiveness of the discriminator, which results in mode collapse—a well-known issue in discriminator-based methods, originally identified in GANs. In such methods, if the discriminator’s performance falls significantly behind the generator (e.g., due to insufficient training or a lack of discriminator expressiveness, which is the case here), the generator will exploit this gap and produce single-mode data
that can easily fool the discriminator. In this case, the generator is falling into a local minima. As the discriminator gradually catches up and learns to identify this data, the generator will discover new data to fool the updated discriminator, falling into another local minimum. This perpetual cycle of chasing leads to continuous oscillations in the training process. We conjecture this is the exact reason for DAC’s instability in this experiments—With the discriminator reduced to a purely linear function, its expressiveness is significantly weakened, allowing the RL policy to exploit it easily, and thus a perpetual chasing emerges. However, our method can remain more effective even with linear score functions, which aligns with our hypothesis in Section 5.1.
7 Conclusion
In this paper, we propose a new imitation learning framework, Score-Matching Imitation Learning (SMILING), that leverages the diffusion score function to learn from expert demonstrations. Different from previous methods, our approach is not formulated from any $f$-divergence or IPM perspective. Instead, we directly match the score function of the learned policy to that of the expert policy. Theoretically, we show that our score function is more expressive than learning a discriminator on $f$-divergence in some sense. Furthermore, we show that our method can achieve first- and second-order bounds in sample complexity, which are instance-dependent and tighter than previous results. In practice, we demonstrate that our method outperforms DAC and IQ-Learn on a set of continuous control tasks, especially on harder tasks.
Appendix A Error Analysis of Objective

### v2

@@ 4 Algorithm
The cost function in Eq. 3 involves the score function $\nabla\log q_{t}(s_{t}\,|\,s)$, which has a closed-form expression for most modern diffusion models including DDPM. Specifically, when the diffusion follows the OU process (described in Section 3.3), the score function takes the form $\nabla\log q_{t}(s_{t}\,|\,s)=(1-e^{-2t})^{-1}(se^{-t}-s_{t})$, which is exactly the noise added to the original sample $s$ over diffusion (recalling that for the OU process, $q_{t}(s_{t}\,|\,s)=\mathcal{N}(se^{-t},(1-e^{-2t})I)$). We denote this noise by $\epsilon$. Then, the cost function (Eq. 3) is equivalent to $c^{(k)}(s):=\mathop{\mathbb{E}}_{t\sim U(T)}\mathop{\mathbb{E}}_{\epsilon\sim\mathcal{N}(0,I)}[\|g^{e}(se^{-t}+\epsilon,t)-\epsilon\|_{2}^{2}-\|g^{(k)}(se^{-t}+\epsilon,t)-\epsilon\|_{2}^{2}].$ This closely resembles the noise prediction in DDPM.
@@ 5 Theoretical Results
    The estimator $g^{e}$ is accurate up to some error $\epsilon_{\rm{score}}$: $\mathop{\mathbb{E}}_{s\sim d^{\pi^{e}}}\mathop{\mathbb{E}}_{t\sim U(T)}\mathop{\mathbb{E}}_{s_{t}\sim q_{t}(\cdot\,|\,s)}[\|g^{e}(s_{t},t)-\nabla_{s_{t}}\log p^{\pi^{e}}_{t}(s_{t})\|_{2}^{2}]\leq\epsilon_{\rm{score}}^{2};$
@@
Item (a) is a standard assumption in diffusion process and score matching (Chen et al., 2023b; a; Lee et al., 2022). When the function class $\mathcal{G}$ (hypothesis space) is finite, $\epsilon^{2}_{\rm{score}}$ typically scales at a rate of $\widetilde{O}\left(\ln(|\mathcal{G}|)/N\right)$ in terms of the number of samples $N$. For an infinite function class, more advanced bounds can be established (e.g., see Theorem 4.3 in Oko et al. (2023)). We emphasize that the discretization error arising from approximating the diffusion SDE using Markov chains is included in $\epsilon_{\rm{score}}$ as well. Item (b) is similar to Item (a) but for the sequence of score function estimators. It can be satisfied by applying Follow-the-Leader (FTL) since the loss functional for $g$ is a square loss. Other no-regret algorithms, such as Follow-the-regularized-leader (FTRL) or online gradient descent, can also be used. Under similar conditions as in Item (a), one typically achieves ${\rm{Regret}}(K)=\widetilde{O}(\sqrt{K})$. Item (c) can be satisfied as long as an efficient RL algorithm is applied at each iteration in 4. The rate of $\epsilon_{\rm{RL}}$ has been well-studied in the RL theory literature for various of MDPs. For instance, in tabular MDPs, one can achieve $\epsilon_{\rm{RL}}=\widetilde{O}(\sqrt{SA/M})$ where $S$ is the number of states, $A$ is the number of actions, and $M$ is the number of RL rollout samples. We emphasize that $\epsilon_{\rm{RL}}$ in this case does not scale with $H$ as it is defined as the upper bound on the averaged return instead of the cumulative return over $H$ steps.
@@
Next, we introduce the regularity assumptions on the diffusion process, which are standard in the literature on diffusion models.
@@
The first condition ensures that the score function behaves well so we can transfer the DS divergence bound to a KL divergence bound. The second condition is needed for the exponential convergence of the forward diffusion process (i.e., the convergence to the standard Gaussian in terms of KL divergence). We note that the second condition is readily satisfied in certain simple cases such as when the state space is bounded. Notably, if we disregard discretization errors, the Lipschitz constant will not appear in our theoretical results as long as it is finite, even if it is arbitrarily large. Similarly, the constant $m$ only appears within logarithmic factors so it can be exponentially large. However, when discretization errors are considered, both the Lipschitz constant and $m$ will appear in $\epsilon_{\rm{score}}$ and ${\rm{Regret}}(T)$ (see, e.g., Chen et al. (2023b; a); Oko et al. (2023)).
Now we are ready to present our main theoretical result. Let $\pi^{(1:K)}$ denote the mixture policy of $\{\pi^{(1)},\dots,\pi^{(K)}\}$. Particularly, $\pi^{(1:K)}$ is executed by first choosing a policy $\pi^{(k)}$ uniformly at random and then executing the chosen policy. For any policy $\pi$, we define the variance of its return as ${\rm{Var}}^{\pi}\coloneqq{\rm{Variance}}(\sum_{h=1}^{H}c^{\star}(s_{h}))$. Below is our main theoretical result with proof in Appendix C.
@@
Under Assumptions 1 and 2, Algorithm 1 achieves these instance-dependent bounds:
@@
where we define total error $\epsilon\coloneqq\begin{cases}\epsilon^{2}_{\rm{score}}+\epsilon_{\rm{RL}}+{\rm{Regret}}(K)/K&\text{if }\pi^{e}\in\Pi,\\ \epsilon^{2}_{\rm{score}}+\epsilon_{\rm{RL}}+{\rm{Regret}}(K)/K+\epsilon_{\rm{mis}}&\text{if }\pi^{e}\not\in\Pi,\end{cases}$ and the misspecification $\epsilon_{\rm{mis}}\coloneqq\min_{\pi\in\Pi}\ell(\pi)$ where we recall that $\ell(\cdot)$ is defined in Eq. 1.
@@
It is known that the second-order bound is tighter and subsumes the first-order bound (Wang et al., 2024c). Importantly, previous results on second-order bounds for sequential decision-making have relied exclusively on the Maximum Likelihood Estimator (MLE) (Wang et al., 2023; 2024c; Foster et al., 2024; Wang et al., 2024b; 2025), which in general is not computationally tractable even for simple exponential family distributions (Pabbaraju et al., 2024). In this work, we show for the first time that a computationally tractable alternative — diffusion score matching — can achieve these instance-dependent bounds in the context of IRL. In particular, our second-order bounds (or the first-order bound) scale with the minimum of ${\rm{Var}}^{\pi^{e}}$ and ${\rm{Var}}^{\pi^{(1:K)}}$ (or minimum of $V^{\pi^{e}}$ and $V^{\pi^{(1:K)}}$). Hence, our bounds are sharper whenever one of the two is small. In contrast, bounds in these prior RL work do not scale with the min of variances associated with the learned policies and the comparator policy $\pi^{e}$. This subtlety is likely due to the difference between the IRL setting and the general RL setting instead of techniques in the analysis.
Improved sample efficiency with respect to the size of expert data. Let’s now focus on sample complexity with respect to expert data, which is typically the most expensive data to collect in IL, to highlight the benefits of our method over previous approaches. For conciseness, we ignore all errors that are purely related to computation. Specifically, we set $\epsilon_{\text{RL}}\rightarrow 0$, as it only depends on the time spent to run the RL procedure, let the number of iterations $K\rightarrow\infty$ so $\text{Regret}(K)/K\rightarrow 0$, and assume $\pi^{e}\in\Pi$. Under these conditions, our bound solely depends on $\epsilon_{\text{score}}$. For simplicity, we further assume a finite score function class $\mathcal{G}$, neglect misspecification errors (i.e., $\nabla\log p^{\pi}_{t}(x)\in\mathcal{G}$ for all $\pi$), and assume the magnitude of the score loss is $O(1)$. Now standard regression analysis (Agarwal et al., 2019; Oko et al., 2023) yields $\epsilon_{\text{score}}=O(\sqrt{\log(|\mathcal{G}|)/N})$ where $N$ is the number of expert samples. Plugging this back into our second-order bound, we obtain
@@
In addition to the sample benefit above, we show our approach can require less expressive function approximators than discriminator-based methods. Please refer to Appendix A for more details.
6 Experiments
We evaluate our method, SMILING, on a set of continuous control tasks from the Deepmind Control Suite (Tassa et al., 2018) and HumanoidBench (Sferrazza et al., 2024). These tasks vary in difficulty, as summarized in Table 1 of Appendix D. While we will compare to multiple baselines including non-adversarial baselines, we focus on testing our hypothesis that score-matching based approach should outperform $f$-divergence-based approach. Additional results and details are in Appendix D.
Baselines. We include Discriminator Actor-Critic (DAC), IQ-Learn, DiffAIL, and Behavior Cloning (BC) as baselines. Our DAC implementation is based on the original design by Kostrikov et al. (2019) that uses JS-divergence GAN-style objective. We use the official implementation of IQ-Learn. We tried several configurations provided in their implementation for the MuJoCo tasks and reported the best one. DiffAIL also follows official implementation. The BC performance is reported as the maximum episode reward obtained across all checkpoints during training. We found MLE outperforms squared loss for BC, so all BC results below use MLE.
RL Solvers. For DMC tasks, we use Soft Actor-Critic (SAC) (Haarnoja et al., 2018) as the RL solver for both DAC and our method, based on the implementation provided by Yarats & Kostrikov (2020). For HumanoidBench tasks, we adopt DreamerV3 (Hafner et al., 2023) using the implementation from the HumanoidBench repository (Sferrazza et al., 2024). Additionally, we maintain a separate “state buffer” that operates the same as the replay buffer but only stores states. This buffer is used to sample states to update the score function in our method and the discriminator in DAC.
Overall, we keep the implementation of SMILING and the baseline DAC as close as possible (e.g., using the same RL solver) so they only differ in objective functions (i.e., SMILING learns score functions while DAC learns discriminators to approximate JS divergence). This allows us to show the benefits of our score-matching approach to the discriminator-based GAN-style ones.
Expert Policy and Demonstrations. We use the aforementioned RL solvers to train expert policies and collect demonstrations by running the expert policy for five episodes. Each task has a fixed 1K-step horizon without early termination, and thus the resulting dataset has 5K steps per task.
6.1 Learning from State-only Demonstrations
Figure 2: Learning curves for learning from state-only data across five random seeds. The x-axis corresponds to the number of environment steps (also the number of policy updates). The y-axis is normalized such that the expert performance is one and the random policy is zero. Our method clearly outperforms all baselines in five tasks out of six.
Figure 2 illustrates the results across six control tasks when the expert demonstrations are state-only. We trained for 0.4M steps on Ball-in-cup-catch and for 3M steps on all other tasks. Our method matches expert performance and outperforms all baselines on five of them. Specifically, our method outperforms all baselines by a large margin on four. For humanoid-sit, while none achieves expert performance, ours converges the fastest. For cheetah-run, although DAC initially shows faster convergence, it becomes unstable soon and oscillates significantly after 2M steps; in contrast, our algorithm remains much more stable and surpasses DAC in 3M steps. Notably, we observed DiffAIL collapses at the beginning of training on all four humanoid tasks due to numerical issues (neural network weights diverging to infinity). Hence, we were unable to report the those results. Also note that, in the plot, we include BC as a reference though it is not directly comparable since it uses state-action demonstrations. Here BC serves as a performance upper bound for Behavioral Cloning from Observation (BCO) (Torabi et al., 2018), which infers the expert’s actions from state-only data before training. Nevertheless, we observe that both our method and DAC consistently outperform BC (and thus, BCO). We provide video demos of SMILING and DAC on the humanoid-crawl and humanoid-pole tasks in Figure 3 to give a clear visual comparison of their performance differences.
6.2 Additional Results
Learning from State-Action Demonstrations. Our method can be easily extended to learning from state-action demonstrations by appending the action vector to the state vector for both training and computing rewards. Hence, we also explore its performance in this setting. The results are consistent with the state-only setting and can be found in Section D.1.
Number of Expert Demonstrations. We also investigate the impact of the number of expert demonstrations on the performance of our method and baselines in order to support our theoretical results in Section 5. The results can be found in Section D.2 and show that our method is clearly more sample-efficient and performs well when the number of expert demonstrations is small.
Expressive of Diffusion Models. In Appendix A, we argued that score matching is more expressive than discriminator-based methods, which may explain why our approach generally outperforms DAC. To support it, we conducted an ablation study on the cheetah-run task where ours converged slower than DAC previously. We removed the activation functions in both the discriminator of DAC and the score function of our method to make them purely linear. We found that DAC’s performance degrades notably while our method remains more effective and outperforms DAC clearly. Detailed results are in Section D.3. This reinforces our claim about the expressiveness of score matching.
7 Conclusion
We propose a new IL framework, Score-Matching Imitation Learning (SMILING), that leverages the diffusion score function to learn from expert demonstrations. Unlike previous methods, SMILING is not formulated from any $f$-divergence or IPM perspective. Instead, it directly matches the score of the learned policy to that of the expert. Theoretically, we show that score function is more expressive than learning a discriminator on $f$-divergence. Additionally, our method achieves first- and second-order bounds in sample complexity, which are instance-dependent and tighter than previous results. In practice, our method outperforms existing algorithms on a set of continuous control tasks.
Acknowledgments
GKS was supported in part by a STTR grant. WS acknowledges support from NSF IIS-2154711, NSF CAREER 2339395, and DARPA LANCER: LeArning Network CybERagents. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
Appendix A Comparison to Discriminator-based (f-divergence and IPM) Methods
@@
Hence, score matching only needs relatively weaker score function class to represent complicated distributions to control certain powerful divergence including KL divergence and Hellinger distance. However, for $f$-divergence-based or IPM-based methods, we need a more expressive discriminator function class to make them serve as a tight lower bound of the divergences. This perhaps explains why in practice for generative models, score matching based approach in general outperforms more traditional GAN based approach. In our experiment, we conduct an ablation study in Section D.3 to compare our approach to a GAN-based baseline under linear function approximation for score functions and discriminators, and we show that SMILING can still learn well even when the score functions are linear with respect to the state $s$.
@@
Appendix B Error Analysis of Objective